# Open-FinLLMs: Open Multimodal Large Language Models for Financial Applications

### v1

@@ 4.2 Performance of Instruction Tuned FinLLaMA-Instruct
Table 9 provides detailed information on the datasets and tasks used for the evaluation of instruction fine-tuned models. We align our evaluations with Fintral’s (Bhatia et al., 2024) settings for consistency and comparability, reporting the average performance across datasets in each task. The evaluation includes 6 tasks and 15 datasets used in the FinBEN paper (Xie et al., 2024a): (1) sentiment analysis (SA) task, using FiQA-SA, FOMC, FPB, and Headlines datasets; (2) named entity recognition (NER) task, using Finer-Ord and NER datasets; (3) number understanding (NU) task, using ConvFinQA and FinQA datasets; (4) text summarization (TS) task, using ECTSUM and EDTSUM datasets; (5) stock movement prediction (SMP) task, using ACL18, BigData22, and CIKM18 datasets; (6) credit scoring (CS) task, using Australia and German datasets.
Table 9: Datasets used for evaluating instruction fine-tuned models. Abbreviations: NU for number understanding, TS for text summarization, SMP for stock movement prediction, and CS for credit scoring.
Task Dataset Test size Metrics License
SA FiQA-SA (Maia et al., 2018) 235 Accuracy Public
FOMC (Shah et al., 2023a) 496 CC BY-NC 4.0
FPB (Malo et al., 2014) 970 CC BY-SA 3.0
Headlines (Sinha and Khandait, 2021) 2,283 CC BY-SA 3.0
NER Finer-Ord (Shah et al., 2023b) 1,080 Entity-F1 CC BY-NC 4.0
NER (Alvarado et al., 2015) 980 CC BY-SA 3.0
NU ConvFinQA (Chen et al., 2022) 1,490 EM Accuracy MIT License
FinQA (Chen et al., 2021) 1,147 MIT License
TS ECTSUM (Mukherjee et al., 2022) 495 Rouge-score Public
EDTSUM (Zhou et al., 2021) 2,000 Public
SMP ACL18 (Xu and Cohen, 2018) 3,720 Accuracy MIT License
BigData22 (Soun et al., 2022) 1,470 Public
CIKM18 (Wu et al., 2018) 1,140 Public
CS Australian (Feng et al., 2024) 690 Accuracy CC BY 4.0
German (Feng et al., 2024) 1,000 CC BY 4.0
@@
Table 10: Performance of FinLLaMA-Instruct and baseline models on benchmark tasks, ranging from 0 to 1. We boldface the best performance in each task and underline the second best-performing models. The results for FinTral are taken from its paper since FinTral has not released its evaluation code or detailed evaluation methodology.
@@
Table 11: Datasets used for multimodal evaluation tasks.
Dataset Test Size Metrics License
MMMU (Yue et al., 2024) 10,500 Accuracy Public
MMMU-Business (Yue et al., 2024) 1,428 Accuracy Public
ChartBench 350 Accuracy our data
TableBench 450 Accuracy our data
Closed Source Gemini-1.5-pro - 49.30 49.80 61.40 58.20
GPT-4o - 55.70 64.30 66.30 66.70
Qwen-VL-MAX - 46.80 39.80 56.00 55.40
Open Source LLaVA-1.5 Vicuna-7B 32.00 26.30 43.40 56.00
LLaVA-1.5 Vicuna-13B 33.60 29.00 49.10 69.10
LLaVA-1.6 Vicuna-7B 32.30 25.60 41.70 42.90
LLaVA-1.6 Vicuna-13B 34.00 29.10 50.30 59.30
Deepseek-VL-7B-Chat DeepSeek-LLM-7B-Base 34.20 28.60 51.40 57.30
Qwen-VL-Chat Qwen-7B 32.00 26.20 52.60 48.20
FinLLaVA FinLLaMA 36.30 30.70 52.90 72.40
@@
Overall, Table 10 shows that FinLLaMA-Instruct outperforms other financial LLMs on 4 out of 6 financial tasks, including surpassing GPT-4 on 3 tasks, underscoring its effectiveness and applicability in the financial domain. FinLLaMA-Instruct also exceeds the performance of all other specialized financial LLMs, including Palmyra-Fin-70B-32, which is significantly larger, in 4 out of 6 financial tasks. In the numerical understanding task, FinLLaMA-Instruct achieves the best performance with an average accuracy score of 0.69, even outperforming GPT-4. This highlights the effectiveness of our instruction tuning using large-scale math reasoning data in enhancing the model’s numeric understanding ability. Furthermore, FinLLaMA-Instruct consistently outperforms Mistral-7B-Instruct on all tasks and surpasses ChatGPT in five out of six tasks, with text summarization being the exception. This underscores its superior performance compared to general LLMs. FinLLaMA-instruct achieves better performance compared with GPT-4 on three key financial analysis tasks, demonstrating the robustness of the FinLLaMA backbone model and the effectiveness of our fine-tuning approach and datasets.
4.3 Performance of Multimodal Extended FinLLaVA
4.3.1 Multimodal Tasks
We evaluate our model on four multimodal understanding tasks as shown in Table 11. The MMMU-Overall dataset, with 10,500 instances, assesses general multimodal capabilities, while the MMMU-Business dataset, with 1,428 instances, evaluates performance in financial domains such as accounting and marketing.
In addition to existing datasets, in this paper, we build two new financial multi-modal evaluation tasks. First, ChartBench tests chart interpretation skills with 350 financially relevant instances selected using the ChartInstructionData. We employed GPT-4o to assign a finance-relevance score, selecting instances with scores of 9 or above (out of 10). These were categorized into seven groups, as detailed in Appendix F.1, with 50 randomly chosen instances per category forming the ChartBench benchmark.
The TableBench assesses multimodal capabilities using tabular images, offering a realistic testbed for handling complex financial data. Our dataset includes 450 questions split between making comparisons and data retrieval tasks, essential for extracting data points and comparing metrics, reflecting key decision-making processes in finance.
4.3.2 Performance
@@
As shown in Table 12, our multimodal model, FinLLaVA, achieves the best performance across all tasks among open-source models with 7B and 13B model sizes. FinLLaVA even outperforms larger models like LLaVA-1.5 and LLaVA-1.6, both of which use Vicuna-13B as backbone. On TableBench, FinLLaVA achieves the best performance and outperforms SOTA commercialized LLMs GPT-4 and Gemini-1.5-pro, which proves the effectiveness of our multimodal extension. These results highlight the robustness and promising performance of FinLLaVA On the TableBench dataset, FinLLaVA not only achieves the best performance but also surpasses state-of-the-art commercialized LLMs like GPT-4 and Gemini-1.5-pro. This success demonstrates the effectiveness of our multimodal extension in enhancing the model’s capability to process and analyze complex financial data. These results highlight its potential for widespread application in the financial domain, offering an efficient solution for interpreting and managing multimodal financial data. Our evaluation differs from FinTral’s in using out-of-domain data, showcasing our model’s robustness and superior performance. We enhance its ability to interpret financial tables by integrating extensive OCR data during alignment. With entirely image-based input, users only need to provide table or chart images, making it convenient for real-world applications like financial reporting and auditing. This simplicity allows financial professionals to efficiently use our model, streamlining their workflow and improving productivity without requiring extensive technical knowledge.
@@ 5 Conclusion
In this paper, we present Open-FinLLMs, an innovative suite of open-source financial language models specifically designed to address the limitations of LLMs in financial applications. Our contributions include FinLLaMA, a foundational model built on the continual pre-training of LLaMA3 8B with an extensive domain-specific dataset, FinLLaMA-Instruct, fine-tuned with diverse instructions for improved instruction-following and conversational capabilities, and FinLLaVA, the multimodal extension capable of handling vision, tabular, and charts. Our comprehensive evaluation, which covers 14 financial tasks with $30$ datasets, one financial trading task, and 4 multimodal tasks, demonstrates that FinLLaMA models outperform the LLaMA3 8B backbone and existing financial LLMs across various scenarios. Notably, FinLLaMA models exhibit comparable performance to GPT-4 in multimodal financial tasks despite their smaller size, highlighting their efficiency and effectiveness. The results underscore the potential of FinLLaMA in advancing Financial AI applications through its robust performance in both language and multimodal financial tasks. By releasing our models and evaluation benchmarks to the public, we aim to foster innovation and collaboration within the financial industry, driving further advancements in financial language modeling and multimodal applications in real-world scenarioes.
Acknowledgements
The authors would like to thank Dr. Chung-chi Chen for providing valuable suggestions that helped improve this work. The authors acknowledge UFIT Research Computing for providing computational resources and support that have contributed to the research results reported in this publication. URL: http://www.rc.ufl.edu. This work is supported by the project JPNP20006 from New Energy and Industrial Technology Development Organization (NEDO). This work has also been partially supported by project MIS 5154714 of the National Recovery and Resilience Plan Greece 2.0 funded by the European Union under the Next Generation EU Program. Xiao-Yang Liu is partly supported by Grant No. IUCRA-CRAFT A24-0104.
Ethical Statement
The development and dissemination of Open FinLLMs by the authors carry full responsibility for any potential violation of rights or arising legal issues. Diligent efforts have been undertaken to ensure the construction of the Open FinLLMs respects privacy and conforms to established ethical guidelines.
This manuscript, inclusive of any associated source codes, models, datasets, and appendices ("Material"), is designated exclusively for academic and educational pursuits. It is crucial to acknowledge that the Material does not provide financial, legal, or investment counsel, nor should it be utilized as a foundation for any form of decision-making.
While the authors have exerted reasonable diligence to verify the accuracy and reliability of the Material, no explicit or implied warranty is extended regarding its completeness or suitability for any specific application. The authors, along with their affiliated entities, absolve themselves of liability for any losses, damages, or other consequences, whether direct or indirect, that may emanate from the employment or reliance upon the Material. It is incumbent upon the user to seek professional consultation for financial, legal, or investment determinations.
By referencing or employing this Material, individuals consent to indemnify, defend, and hold the authors, along with any affiliated organizations or persons, harmless against any claims or damages that may arise from such utilization.
Appendix A Contributions
Science Leadership: Qianqian Xie, Min Peng, Sophia Ananiadou, Alejandro Lopez-Lira, Yanzhao Lai, Benyou Wang, Xiao-yang Liu, Jordan W. Suchow, Honghai Yu, Jianyun Nie, Jimin Huang.
Contributors: Dong Li, Mengxi Xiao, Zihao Jiang, Ruoyu Xiang, Xiao Zhang, Zhengyu Chen, Yueru He, Weiguang Han, Yuzhe Yang, Shunian Chen, Yifei Zhang, Lihang Shen, Daniel Kim, Zhiwei Liu, Zheheng Luo, Yangyang Yu, Yupeng Cao, Zhiyang Deng, Zhiyuan Yao, Haohang Li, Duanyu Peng, Yongfu Dai, VijayaSai Somasundaram, Peng Lu, Yilun Zhao, Yitao Long, Guojun Xiong, Kaleb Smith
@@ B.1 Dataset Details
Our continual pre-training corpus is designed to ensure comprehensive coverage of financial knowledge by integrating a diverse range of data sources. This appendix provides a detailed overview of each data source, including the time range covered:
  • Financial Papers: This subset includes 4 billion tokens extracted from academic papers and research articles, offering a strong foundation in financial concepts and theories. The papers span a period from 2000 to 2023, covering a wide range of topics such as market analysis, financial modeling, and economic theory. These documents are sourced from SSRN [^6] https://www.ssrn.com/index.cfm/en/ and open-source conference proceedings, providing in-depth insights into both foundational and cutting-edge financial research.
@@ C.1 Data Sources
  • ChanceFocus/FLUPE (Xie et al., 2023b): The FLUPE dataset, is instrumental in improving financial natural language processing capabilities. It includes tasks such as financial sentiment analysis, news headline classification, and named entity recognition (NER). These tasks involve analyzing financial texts to identify sentiment, classify financial headlines, and recognize entities within financial documents. By providing diverse examples, FLUPE helps models refine their understanding of financial language and context, supporting improved task performance across various financial NLP applications.
  • FinGPT/Fingpt-finred (Liu et al., 2023): This dataset focuses on financial report and document comprehension, with approximately 32.67k examples. It is designed to enhance the model’s ability to interpret complex financial documents. The dataset enables models to improve their analytical skills and decision-making abilities based on comprehensive document analysis.
@@ D.1 Table
Our data is selected from the Fintabnet and Marketing categories of SynthTabNet. The tables are extracted from reports and other PDF documents, featuring real financial data, marketing data, and diverse styling appearances. Additionally, this dataset includes parsed bounding boxes, which enable us to reconstruct the table structure using the prompt as following content. This method is more accurate and informative than generating descriptions or instructions directly from images.
@@
- Generate a descriptive question that encourages a deep dive into the table’s displayed data and its organizational framework. - Avoid using HTML tags in your response. - Ensure your response contains numerical data. - Provide detailed information about the table’s content and structure, highlighting specific values from the table (excluding headers) - Provide spatial information about the cells in this table without including any bounding box information. For example, you can describe the column or row layout. Tasks for SFT Data:
@@
Additionally, some table images have high resolutions, which can cause individual cell information to become blurry after processing with an image processor, negatively impacting the quality of OCR. Therefore, when selecting data, we have limited the rows and columns of the tables to a maximum of $10*10$.
In the SFT stage, we designed seven specific tasks in Appendix F.1 to enable the model to better understand and perform common tasks in the financial domain. In the prompts, tasks are randomly selected from these seven categories. Figure 2 presents an example table image, demonstrating how we align and generate SFT data for specific table tasks.
@@
D.2 Chart
Our chart dataset is derived from Unichart, Chart2Text, ChartQA and features a diverse array of charts displaying real financial data, marketing trends, and various visual styles. We focus on numerical content and financial charts, offering a robust resource for quantitative analysis. Our prompts are shown below and an example question-answer pair is shown in Figure 4.
In the SFT stage, we developed a set of seven specialized tasks (as shown in Appendix F.1) to enhance the model’s ability to comprehend and perform essential operations within the financial domain. These tasks are specifically designed to focus on the interpretation of financial data, conducting detailed analyses, and extracting meaningful insights from chart information. Figure 4 presents an example chart image, demonstrating how we align and generate SFT data for specific chart tasks.
@@ Appendix F Experiments
F.1 Multi-Modal Task Categories

### v2

@@ 3.2 Performance of FinLLaMA-Instruct
Table 10 summarizes the datasets and tasks used to evaluate instruction fine-tuned models, aligned with FinTral’s (Bhatia et al., 2024) settings for consistency. The evaluation spans 6 tasks and 15 datasets from FinBEN (Xie et al., 2024a): (1) sentiment analysis (FiQA-SA, FOMC, FPB, Headlines); (2) named entity recognition (Finer-Ord, NER); (3) number understanding (ConvFinQA, FinQA); (4) text summarization (ECTSUM, EDTSUM); (5) stock movement prediction (ACL18, BigData22, CIKM18); and (6) credit scoring (Australia, German).
Closed Source Gemini-1.5-pro - 49.3 49.8 61.4 58.2
GPT-4o - 55.7 64.3 66.3 66.7
Qwen-VL-MAX - 46.8 39.8 56.0 55.4
Open Source LLaVA-1.5 Vicuna-7B 32.0 26.3 43.4 56.0
LLaVA-1.5 Vicuna-13B 33.6 29.0 49.1 69.1
LLaVA-1.6 Vicuna-7B 32.3 25.6 41.7 42.9
LLaVA-1.6 Vicuna-13B 34.0 29.1 50.3 59.3
Deepseek-VL-7B-Chat DeepSeek-LLM-7B-Base 34.2 28.6 51.4 57.3
Qwen-VL-Chat Qwen-7B 32.0 26.2 52.6 48.2
FinLLaVA FinLLaMA 36.3 30.7 52.9 72.4
@@
Table 7: Performance of FinLLaMA-Instruct and baselines on benchmark tasks (0-1 range). Best performances are bolded, and second-best are underlined. FinTral results are cited from its paper due to unavailable evaluation code or methodology.
@@
Overall, Table 7 shows that FinLLaMA-Instruct shows that FinLLaMA-Instruct outperforms other financial LLMs on 4 out of 6 financial tasks, including surpassing GPT-4 on 3 tasks, underscoring its effectiveness and applicability in the financial domain. FinLLaMA-Instruct also exceeds the performance of all other specialized financial LLMs, including Palmyra-Fin-70B-32, which is significantly larger, in 4 out of 6 financial tasks. In the numerical understanding task, FinLLaMA-Instruct achieves the best performance with an average accuracy score of 0.69, even outperforming GPT-4. This highlights the effectiveness of our instruction tuning using large-scale math reasoning data in enhancing the model’s numeric understanding ability. Furthermore, FinLLaMA-Instruct consistently outperforms Mistral-7B-Instruct on all tasks and surpasses ChatGPT in five out of six tasks, with text summarization being the exception. This underscores its superior performance compared to general LLMs. FinLLaMA-instruct achieves better performance compared with GPT-4 on three key financial analysis tasks, demonstrating the robustness of the FinLLaMA backbone model and the effectiveness of our fine-tuning approach and datasets.
3.3 Performance of FinLLaVA
Multimodal Tasks We evaluate our model on four multimodal understanding tasks (Table 11). The MMMU-Overall dataset (10,500 instances) assesses general multimodal capabilities, while MMMU-Business (1,428 instances) focuses on financial domains such as accounting and marketing. In addition to existing datasets, we introduce two new financial multimodal benchmarks. ChartBench tests chart interpretation skills with 350 finance-related instances selected from ChartInstructionData. We used GPT-4o to assign finance-relevance scores and selected samples scoring 9 or above (out of 10), categorized into seven groups (Appendix F.3), with 50 randomly sampled instances per category. TableBench evaluates multimodal reasoning over tabular images, providing a realistic setting for complex financial data interpretation. It contains 450 questions across comparison and data retrieval tasks, reflecting key decision-making processes in finance.
@@ 3.3 Performance of FinLLaVA
Performance As shown in Table 6, our multimodal model, FinLLaVA, achieves the best performance across all tasks among open-source 7B and 13B models, even surpassing larger models like LLaVA-1.5 and LLaVA-1.6 (Vicuna-13B). On TableBench, FinLLaVA not only outperforms open-source baselines but also surpasses commercial state-of-the-art models such as GPT-4 and Gemini-1.5-pro, demonstrating the effectiveness of our multimodal extension for complex financial data. These results highlight FinLLaVA’s robustness and generalization, especially given that our evaluation uses out-of-domain data, unlike FinTral. By incorporating extensive OCR data during alignment, FinLLaVA improves its ability to interpret financial tables. With fully image-based inputs, users can directly provide table or chart images, making it practical for real-world applications like financial reporting and auditing. This simplicity enables financial professionals to streamline workflows and enhance productivity without technical expertise. Overall, FinLLaVA achieves state-of-the-art performance across multimodal financial tasks, surpassing both open-source and commercial models, with strong generalization and practical utility for real-world applications like financial reporting and auditing.
@@ 4 Conclusion
In this paper, we present Open-FinLLMs, an innovative suite of open-source financial language models specifically designed to address the limitations of LLMs in financial applications. Our contributions include FinLLaMA, a foundational model built on the continual pre-training of LLaMA3 8B with an extensive multimodal domain-specific datasets, FinLLaMA-Instruct, fine-tuned with diverse instructions for improved instruction-following and conversational capabilities, and FinLLaVA, the multimodal extension capable of handling vision, tabular, and charts. Our comprehensive evaluation, which covers 14 financial tasks with $30$ datasets and 4 multimodal tasks, demonstrates that Open-FinLLMs models outperform existing financial LLMs and even better performance to GPT-4 and GPT-4o despite their smaller size, on financial NLP, reasoning, decision-making and multimodal tasks in zero-shot, few-shot, and supervised fine-tuning settings. The results underscore the potential of Open-FinLLMs in advancing Financial AI applications through their robust performance in various financial tasks.
Appendix A Related Work
A.1 Financial Large Language Models
Recently, several financial LLMs have been developed to address domain-specific challenges. BloombergGPT (Wu et al., 2023a) pioneered financial LLMs with pretraining on 363 billion financial tokens; however, it remains closed-source, limiting its accessibility. On the open-source front, PIXIU (Xie et al., 2023b) and FinGPT (Liu et al., 2023, 2024c; Yang et al., 2023) fine-tuned LLaMA models with hundreds of thousands of financial instructions, focusing primarily on text-based tasks. FinTral (Bhatia et al., 2024) introduced multimodal capabilities for tabular data, but limited by the small size of its domain-specific data and its evaluation scope. Fin-o1 Qian et al. (2025), developed through instruction tuning on reasoning-enhanced data, demonstrates that domain-enhanced reasoning abilities significantly improve the model’s performance in reasoning tasks.
A.2 Domain Specialization of Large Language Models
Domain-specific LLMs are developed using three main strategies: pre-training from scratch (PT), continued pre-training (CPT), and instruction fine-tuning (SFT) (Wu et al., 2024). CPT adapts existing LLMs by further training on domain data, like Code LLaMA (Roziere et al., 2023), which improves code generation. SFT tailors models for specific tasks with domain instructions, requiring less data and compute, as demonstrated by PIXIU (Xie et al., 2023b) and FinGPT (Liu et al., 2024c).
@@ B.1 Dataset Details
Our continual pre-training corpus is designed to ensure comprehensive coverage of financial knowledge by integrating a diverse range of data sources. This appendix provides a detailed overview of each data source:
  • Financial Papers: This subset includes 4 billion tokens extracted from academic papers and research articles, offering a strong foundation in financial concepts and theories. The papers span a period from 2000 to 2023. These documents are sourced from SSRN and open-source conference proceedings, providing in-depth insights into both foundational and cutting-edge financial research.
@@ C.1 Data Sources
  • ChanceFocus/FLUPE (Xie et al., 2023b): The FLUPE dataset, is instrumental in improving financial natural language processing capabilities. It includes tasks such as financial sentiment analysis, news headline classification, and named entity recognition (NER). These tasks involve analyzing financial texts to identify sentiment, classify financial headlines, and recognize entities within financial documents. By providing diverse examples, FLUPE helps models refine their understanding of financial language and context, supporting improved task performance across various financial NLP applications. FinGPT/Fingpt-finred (Liu et al., 2023): This dataset focuses on financial report and document comprehension, with approximately 32.67k examples. It is designed to enhance the model’s ability to interpret complex financial documents. The dataset enables models to improve their analytical skills and decision-making abilities based on comprehensive document analysis.
@@ D.1 Table
Our table data is sourced from the Fintabnet and Marketing categories of SynthTabNet, featuring real financial and marketing tables with diverse layouts. Each table includes parsed bounding boxes, allowing us to reconstruct structure-aware prompts, which are more accurate than image-based descriptions.
To ensure OCR quality, we limit tables to a maximum size of $10\times 10$ to avoid resolution-related cell blurring. In the SFT stage, we design seven financial-specific tasks (Appendix F.3), randomly sampled in the prompts. Figure 2 shows an example table and how we align and generate SFT data.
D.2 Chart
Our chart dataset is derived from Unichart, Chart2Text, and ChartQA, covering real financial data, marketing trends, and varied visual styles. We focus on numerical and financial charts to support robust quantitative analysis. An example question-answer pair is shown in Figure 3.
During SFT, we develop seven chart-specific tasks (Appendix F.3) to enhance the model’s ability to interpret financial charts and extract insights. Figure 4 illustrates an example chart used in SFT data generation.
@@ D.2 Chart
- Generate a descriptive question that encourages a deep dive into the table’s displayed data and its organizational framework. - Avoid using HTML tags in your response. - Ensure your response contains numerical data. - Provide detailed information about the table’s content and structure, highlighting specific values from the table (excluding headers). - Provide spatial information about the cells in this table without including any bounding box information. For example, you can describe the column or row layout. Tasks for SFT Data:
@@ Appendix F Experiments
F.1 Datasets used for different experiments
Details of the dataset used for evaluation are listed in Table  8, 9, 10, 11.
Table 8: Datasets used for zero-shot evaluation tasks. The dataset with * is newly created in this paper. Abbreviations: EM Accuracy means exact match accuracy. MCC means Matthews correlation coefficient.
Task Dataset Test Size Metrics License
Sentiment analysis TSA (Cortis et al., 2017) 561 F1 CC BY-NC-SA 4.0
Classification FOMC (Shah et al., 2023a) 496 F1 CC BY-NC 4.0
Classification FinArg-AUC (Chen et al., 2023) 969 F1 CC BY-NC-SA 4.0
Classification MA (Yang et al., 2020) 500 F1 Public
Classification SC (Mariko et al., 2020) 8,630 Entity F1 CC BY 4.0
Math *MC 50 F1 Public
Math KnowledgeMath (Zhao et al., 2023a) 1000 EM Accuracy MIT License
Math DocMath-Eval (Zhao et al., 2023b) 3200 EM Accuracy MIT License
Credit scoring German (Hofmann, 1994) 1,000 MCC CC BY 4.0
Credit scoring Australian (Quinlan, ) 690 MCC CC BY 4.0
Credit scoring LendingClub (Feng et al., 2024) 2,690 MCC CC0 1.0
Fraud detection ccf (Feng et al., 2024) 2,278 MCC DbCL v1.0
Fraud detection ccfraud (Feng et al., 2024) 2,097 MCC Public
Distress polish (Feng et al., 2024) 1,736 MCC CC BY 4.0
Distress taiwan (Feng et al., 2024) 1,364 MCC CC BY 4.0
Claim analysis ProtoSeguro (Feng et al., 2024) 2,381 MCC Public
Claim analysis travelinsurance (Feng et al., 2024) 3,800 MCC ODbL v1.0
Question answering ConvFinQA (Chen et al., 2022) 1,490 EM Accuracy MIT License
Table 9: Datasets used for few-shot evaluation tasks.
Task Dataset Test Size Metrics License
Sentiment analysis FPB (Malo et al., 2014) 970 F1 CC BY-SA 3.0
Sentiment analysis FiQA-SA (Maia et al., 2018) 235 F1 Public
Named entity recognition NER (Alvarado et al., 2015) 980 Entity F1 CC BY-SA 3.0
Table 10: Datasets used for evaluating instruction fine-tuned models.
Task Dataset Test size Metrics License
Sentiment analysis FiQA-SA (Maia et al., 2018) 235 Accuracy Public
FOMC (Shah et al., 2023a) 496 CC BY-NC 4.0
FPB (Malo et al., 2014) 970 CC BY-SA 3.0
Headlines (Sinha and Khandait, 2021) 2,283 CC BY-SA 3.0
Named entity recognition Finer-Ord (Shah et al., 2023b) 1,080 Entity-F1 CC BY-NC 4.0
NER (Alvarado et al., 2015) 980 CC BY-SA 3.0
Number understanding ConvFinQA (Chen et al., 2022) 1,490 EM Accuracy MIT License
FinQA (Chen et al., 2021) 1,147 MIT License
Text summarization ECTSUM (Mukherjee et al., 2022) 495 Rouge-score Public
EDTSUM (Zhou et al., 2021) 2,000 Public
Stock movement prediction ACL18 (Xu and Cohen, 2018) 3.720 Accuracy MIT License
BigData22 (Soun et al., 2022) 1,470 Public
CIKM18 (Wu et al., 2018) 1,140 Public
Credit scoring Australian (Quinlan, ) 690 Accuracy CC BY 4.0
German (Hofmann, 1994) 1,000 CC BY 4.0
Table 11: Datasets used for multimodal evaluation tasks.
Dataset Test Size Metrics License
MMMU (Yue et al., 2024) 10,500 Accuracy Public
MMMU-Business (Yue et al., 2024) 1,428 Accuracy Public
ChartBench 350 Accuracy our data
TableBench 450 Accuracy our data
F.2 Details of Agent evaluation results
Results of different models in agent evaluation setting is lised in F.2.
Table 12: Trading performance of FinLLaMA and baseline models.
Metric Asset Buy & Hold LLaMA3.1-8B LLaMA3-8B FinLLaMA
Cumulative Return TSLA -0.1633 -0.1829 -0.1552 0.5573
COIN -0.0562 -0.3515 -0.0319 0.1743
GOOG -0.0562 0.1651 0.1631 0.1098
NIO -0.3530 -0.2677 -0.3526 0.4645
Overall -0.1571 -0.1592 -0.0942 0.3265
Sharpe Ratio TSLA -0.4769 -0.5557 -0.8486 2.4532
COIN -0.1041 -0.7911 -0.0666 0.5778
GOOG -0.1041 0.9464 1.0128 0.6927
NIO -0.8678 -0.6951 -1.0313 1.9113
Overall -0.3823 -0.2739 -0.2334 1.4088
Normalized Sharpe Ratio TSLA 42.05 40.72 35.86 90.89
COIN 48.27 36.82 48.89 59.63
GOOG 48.27 82.44 83.88 61.55
NIO 35.54 38.41 32.81 81.86
Overall 43.63 45.43 45.92 73.48
Annual Volatility TSLA 0.7015 0.6741 0.3748 0.4654
COIN 1.1047 0.9103 0.9823 0.6181
GOOG 1.1047 0.3574 0.3300 0.3249
NIO 0.8332 0.7891 0.7005 0.4979
Overall 0.9360 0.6827 0.5969 0.4766
Max Drawdown TSLA 0.5532 0.5248 0.2662 0.1883
COIN 0.6019 0.5449 0.4056 0.4142
GOOG 0.6019 0.1983 0.2092 0.1741
NIO 0.4498 0.4213 0.5172 0.3008
Overall 0.5517 0.4223 0.3496 0.2693
F.3 Multimodal Task Categories